# Relational Dissonance in Human-AI Interactions: The Case of Knowledge Work

### v2

@@ Abstract.
$c$s: Human-centered computing Natural language interfacesccs: Human-centered computing HCI theory, concepts and models
Figure 1. Core components of the relational-dissonance framework, illustrated on data collected from knowledge workers. :
@@ 1. Introduction
Defining relational dissonance. Relational dissonance describes the divergence between (1) an individual’s ontological and relational framing of AI systems and (2) the relational dynamics that emerge as their interaction with these systems unfolds. In our conception, relational dissonance captures the difficulty of keeping stable answers to two questions about AI: “What is AI?” (ontological ambiguity) and “What is AI to me?” (relational ambiguity). The resulting dynamics can lead to a lack of alignment between individuals’ articulated accounts (explicit dimension) and enacted relationships (implicit dimension), which remains mostly unnoticed during an individual’s interaction with a system. We understand relational dynamics as observable qualities in human-AI interactions, including the roles enacted and the social tone of the exchange. Building on recent HCI research (Kegel et al., 2025; Fu et al., 2025), we use the term relational to differentiate socially complex interactions from purely transactional exchanges. Our focus lies on the live, momentary dynamics of a person interacting with what Turkle et al. (Turkle et al., 2006) call a relational artifact, i.e., an entity that invites partnership-like responses without having subjectivity or needs. To the best of our knowledge, our definition of relational dissonance creates a novel theoretical construct specific to human-AI interactions, completely distinct from a prior narrow appearance of the term in adolescent psychology (Bond et al., 2014).
@@ 2. Related Work: Understanding AI Ontology and Relationality in Human-AI Interactions
Researchers have tried to comprehend AI systems from various angles. At the ontological level, research concentrates on understanding with what, or with whom, people are interacting (Etzrodt, 2021). This body of work largely emphasizes the instrumental use of AI, foregrounding efficiency and other benefits, often with limited attention to the individual. Within the managerial literature, most studies examine the impact of AI technologies on a range of workplace outcomes (Budhwar et al., 2023; Pereira et al., 2023), highlighting their positive effects on individual performance, satisfaction, commitment, and employee experiences (Malik et al., 2022), which, in turn, shape firm-level outcomes (Lee et al., 2023; Malik et al., 2022).
@@ 3.1. Study Design and Workshop Protocol
Figure 2. Workshop structure. Progressing from individual to collective activities, our workshops included (a) a pre-session survey capturing participants’ initial perceptions of AI systems and their use in qualitative research; (b) individual hands-on interaction sessions with AI systems for qualitative-data-analysis tasks; (c) a peer observation activity where participants reviewed and reflected on one another’s AI interactions; (d) an image selection and description exercise to elicit metaphorical representations of AI; (e) semi-structured group discussion exploring collective experiences and sense-making; and (f) a collaborative worksheet activity documenting shared insights. :
@@ 4. Findings
Data notes. Two participants did not fill out the pre-session survey (P2.7 and P2.8), three did so only partially (P1.6,P2.5, P3.6), and one did not specify their gender or age (P3.2). For one participant (P1.4), we were unable to retrieve the chatlog. :
@@ 4.2. Oscillating Relational Configurations
The process of negotiation and mutual adaptation, illustrated in the example, repositions the system’s output not as a static product of technology but as a direct outcome of the relational configurations between humans and AI. Our attempts to impose discrete phases on fundamentally continuous relational adjustments reveal the limitations of segmented approaches. The most important finding of this analysis is not which categories individuals or AI systems belong to, but rather that they do not remain in any category long enough for such classifications to holistically capture the human-AI relational space. This highlights the need to develop approaches that are better suited to capture the relational dynamics that emerge when humans interact with anthropomorphic conversational agents. Relational dissonance takes one step toward this goal, and we now proceed to show how it manifests in a social setting.
Figure 4. Visual metaphors selected by the knowledge workers who participated in our study. Images are scaled by the number of participants who selected them, and participants who selected an image are represented by badges and colored by their dominant relational configuration (see Figure 1). Four participants selected two visual metaphors instead of just one metaphor (P3.1, P3.3, P3.4, P3.7). For the full set of images presented to participants during the image-selection exercise, see Figure 5 in the Appendix.:
@@ 5. Discussion
Our workshops with qualitative researchers revealed a fundamental tension in human-AI interactions: a discrepancy between people’s explicit framing of AI systems and the relational dynamics that manifest during their actual engagement with these systems. We term this divergence relational dissonance, shaping a novel analytical construct to understand the implicit dynamics that ultimately determine the processes and outcomes of knowledge work. In this section, we first clarify the foundational role of relational dissonance in human-AI interactions and its implications for the instrumental use of AI in knowledge work. We then explore the concept of relational transparency to address possible methodological, epistemic, and personal risks resulting from human-AI relational dynamics.
@@
The Entanglement of Relational and Analytical Work. Contemporary discourse on AI in knowledge work is dominated by ‘tool use’ narratives. Industry leaders urge professionals to ‘master this new tool’ (Webb, 2025), universities offer prompt-engineering courses or guides on ‘how to use AI’ (web, 2025). These framings position AI as a sophisticated but controllable instrument that can be integrated into existing workflows with proper training and methods. The emphasis on prompting strategies, best practices, and optimization techniques delivers the message that successful human-AI interaction is primarily a matter of technical proficiency.
@@ 6. Limitations
Conceptual Limitations. While this study conceptually distinguishes relational dissonance from related constructs such as cognitive dissonance and the intention-behavior gap, our empirical evidence is mostly qualitative, based on discrepancies between explicit accounts and implicit dynamics identified across our multiple data sources. Future research could focus on developing more direct measures to quantitatively substantiate relational dissonance. We acknowledge that, as researchers already invested in a relational perspective, our analytical lens inevitably shaped our interpretation. Although our findings are based on specific participant accounts and chatlogs, it is important to view our framework as just one interpretive lens among many. Moreover, as relational dissonance partially operates outside conscious awareness, our methodology captures its observable manifestations, rather than the complete psychological process of relational dissonance. The value of this construct lies in illuminating the continuous micro-negotiations between explicit framings and implicit dynamics, which cannot be fully captured by HCI taxonomies that assume stable roles.
@@
Empirical Limitations. Our findings emerge from a specific population of qualitative researchers operating within the professional norms of knowledge production, agency, and methodological reflexivity. We cannot claim that the relational dynamics we document are universal or representative of the general population. The tensions between tool-use framing and relational engagement might be especially pronounced in our study population, which has professional stakes in interpreting qualitative data. As entirely different modes of relational engagement may arise during everyday interactions with AI systems, generalizing our findings to a general audience is not straightforward. Therefore, our empirical contribution should be understood as documenting how relational dissonance manifests in knowledge-work contexts where professional integrity and epistemic authority are at stake, rather than identifying universal dynamics of human-AI interaction.
@@ Appendix C Study Details
In this appendix, we provide key details related to the setup of our study, sharing (1) the mapping between the human-readable participant identifiers used in the paper and the random identifiers used in the study itself (Section C.1), (2) the survey questions and item scales used in the pre-session survey (Section C.2), (3) the precise wording of the fictional qualitative-research task to be addressed by participants through human-AI interactions (Section C.3), (4) the precise versions of the AI systems used by our participants (Section C.4), (5) the images presented during the metaphor-elicitation exercise (Section C.5), and (6) the structure underlying the group discussions (Section C.6).
@@ C.2. Survey Questions and Item Scales
In the following, we reproduce the questions participants answered in our pre-session survey, organized by their variable group in Fig. 3, along with the scales on which they could provide their answer (see Section 3.1). In addition to these questions, we also asked participants to self-report their age, nationality, and gender as basic demographic information. All questions were optional. An anonymized CSV file of all answers (with nationalities censored to protect privacy) is provided in the supplementary material.
@@ C.5. Images Provided for Selection as Visual Metaphors
Figure 5. Visual metaphors used in the image-selection exercise. :

### v3

# arXiv:2509.15836v3[cs.HC] 27 Feb 2026

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791180isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Natural language interfacesccs: Human-centered computing HCI theory, concepts and models
Figure 1. Core components of the relational-dissonance framework, illustrated on data collected from knowledge workers.
Three-column figure illustrating the relational-dissonance framework. In the left column, we show five typical relational configurations in knowledge work, along with their net initiative energy: Director (human gives unidirectional commands to AI), Trainer (human guides with bidirectional feedback), Partner (equal bidirectional exchange), Student (AI guides human), and Consumer (human passively receives from AI). In the middle and right column, we illustrate the relational-dissonance concept, showing misalignment between relational expectations and relational experiences based on survey responses and chat excerpts from two study participants (P1.7 middle, P3.2 right) showing dialogue exchanges with AI systems. For both participants, the survey responses indicate Director-type expectations, but the actual chat interactions demonstrate multiple shifts toward configurations with less human control.
@@ 1. Introduction
Defining relational dissonance. Relational dissonance describes the divergence between (1) an individual’s ontological and relational framing of AI systems and (2) the relational dynamics that emerge as their interaction with these systems unfolds.
In our conception, relational dissonance captures the difficulty of keeping stable answers to two questions about AI: “What is AI?” (ontological ambiguity) and “What is AI to me?” (relational ambiguity). The resulting dynamics can lead to a lack of alignment between individuals’ articulated accounts (explicit dimension) and enacted relationships (implicit dimension), which remains mostly unnoticed during an individual’s interaction with a system. We understand relational dynamics as observable qualities in human-AI interactions, including the roles enacted and the social tone of the exchange. Building on recent HCI research (Kegel et al., 2025; Fu et al., 2025), we use the term relational to differentiate socially complex interactions from purely transactional exchanges. Our focus lies on the live, momentary dynamics of a person interacting with what Turkle et al. (Turkle et al., 2006) call a relational artifact, i.e., an entity that invites partnership-like responses without having subjectivity or needs. To the best of our knowledge, our definition of relational dissonance creates a novel theoretical construct specific to human-AI interactions, completely distinct from a prior narrow appearance of the term in adolescent psychology (Bond et al., 2014).
@@ 2. Related Work: Understanding AI Ontology and Relationality in Human-AI Interactions
Researchers have tried to comprehend AI systems from various angles. At the ontological level, research concentrates on understanding with what, or with whom, people are interacting (Etzrodt, 2021). This body of work largely emphasizes the instrumental use of AI, foregrounding efficiency and other benefits, often with limited attention to the individual. Within the managerial literature, most studies examine the impact of AI technologies on a range of workplace outcomes (Budhwar et al., 2023; Pereira et al., 2023), highlighting their positive effects on individual performance, satisfaction, commitment, and employee experiences (Malik et al., 2023), which, in turn, shape firm-level outcomes (Lee et al., 2023; Malik et al., 2022).
@@ 3.1. Study Design and Workshop Protocol
Figure 2. Workshop structure. Progressing from individual to collective activities, our workshops included (a) a pre-session survey capturing participants’ initial perceptions of AI systems and their use in qualitative research; (b) individual hands-on interaction sessions with AI systems for qualitative-data-analysis tasks; (c) a peer observation activity where participants reviewed and reflected on one another’s AI interactions; (d) an image selection and description exercise to elicit metaphorical representations of AI; (e) semi-structured group discussion exploring collective experiences and sense-making; and (f) a collaborative worksheet activity documenting shared insights.
Six-panel diagram illustrating our workshop methodology flow. Panel (a) shows a pre-session survey form with checkboxes. Panel (b) depicts a participant at a computer interacting with their assigned AI system. Panel (c) illustrates peer observation activity, where participants view one another’s screens. Panel (d) illustrates an image selection exercise in which participants choose from visual metaphors. Panel (e) displays a semi-structured group discussion with participants seated in a circle. Panel (f) shows a collaborative worksheet activity with participants at a table using post-it notes. Red arrows connect the sequential phases from individual to collective activities.
@@ 4. Findings
Data notes. Two participants did not fill out the pre-session survey (P2.7 and P2.8), three did so only partially (P1.6,P2.5, P3.6), and one did not specify their gender or age (P3.2). For one participant (P1.4), we were unable to retrieve the chatlog.
Two-panel figure displaying a subset of the data sources underlying our study: (a) Pre-session survey responses showing scores across 4 variables for 22 participants, with responses ranging from negative (0) to positive (1) on performance expectancy, effort expectancy, attitudes toward use, and satisfaction. (b) Horizontal bar chart showing chatlog token counts (0-5000) for each participant’s interaction with AI systems. Participant IDs and chatlogs are colored by their dominant relational configuration.
@@ 4.2. Oscillating Relational Configurations
The process of negotiation and mutual adaptation, illustrated in the example, repositions the system’s output not as a static product of technology but as a direct outcome of the relational configurations between humans and AI. Our attempts to impose discrete phases on fundamentally continuous relational adjustments reveal the limitations of segmented approaches. The most important finding of this analysis is not which categories individuals or AI systems belong to, but rather that they do not remain in any category long enough for such classifications to holistically capture the human-AI relational space. This highlights the need to develop approaches that are better suited to capture the relational dynamics that emerge when humans interact with anthropomorphic conversational agents. Relational dissonance takes one step toward this goal. We now proceed to show how it manifests in a social setting.
Figure 4. Visual metaphors selected by the knowledge workers who participated in our study. Images are scaled by the number of participants who selected them, and participants who selected an image are represented by badges and colored by their dominant relational configuration (see Figure 1). Four participants selected two visual metaphors instead of just one metaphor (P3.1, P3.3, P3.4, P3.7). For the full set of images presented to participants during the image-selection exercise, see Figure 5 in the Appendix.
Gallery presenting the images selected by the participants in our study. The figure shows two rows of images that are scaled by the number of participants who selected them. Each image is annotated with the identifiers of the participants who selected it, who are represented by badges in the color of their dominant relational configuration.
@@ 5. Discussion
Our workshops with qualitative researchers revealed a fundamental tension in human-AI interactions: a discrepancy between people’s explicit framing of AI systems and the relational dynamics that manifest during their actual engagement with these systems. We term this divergence relational dissonance, shaping a novel analytical construct to understand the implicit dynamics that ultimately determine the processes and outcomes of knowledge work. In this section, we first clarify the role of relational dissonance in human-AI interactions and its implications for the instrumental use of AI in knowledge work. We then explore the concept of relational transparency to address possible methodological, epistemic, and personal risks resulting from human-AI relational dynamics.
@@
The Entanglement of Relational and Analytical Work. Contemporary discourse on AI in knowledge work is dominated by ‘tool use’ narratives. Industry leaders urge professionals to ‘master this new tool’ (Webb, 2025), universities offer prompt-engineering courses or guides on ‘how to use AI’ (33). These framings position AI as a sophisticated but controllable instrument that can be integrated into existing workflows with proper training and methods. The emphasis on prompting strategies, best practices, and optimization techniques delivers the message that successful human-AI interaction is primarily a matter of technical proficiency.
@@
Qualitative research practice has long been characterized by scholars’ continuous examination of their own interpretive influence on their data. When using AI systems in their work, qualitative researchers must maintain this reflexivity not only toward themselves but also toward AI systems as contributors to the interpretive process. However, the relational complexities within the human-AI interface make the continuous reflexivity required by existing qualitative frameworks harder to enact, with implications for the development of future qualitative methods.
@@ 6. Limitations
Conceptual Limitations. While this study conceptually distinguishes relational dissonance from related constructs such as cognitive dissonance and the intention-behavior gap, our empirical evidence is mostly qualitative, based on discrepancies between explicit accounts and implicit dynamics identified across our multiple data sources. Future research could focus on developing more direct measures to quantitatively substantiate relational dissonance. For example, specific conversational moves—such as praise, greetings, evaluative feedback, or pushback—could form the basis of a coding scheme for creating observable markers of shifts in relational stance, and divergences between people’s explicit articulations or initial exchanges with an AI system and what their interaction traces subsequently reveal could be operationalized as indicators of the resulting structural tension. We further acknowledge that, as researchers already invested in a relational perspective, our analytical lens inevitably shaped our interpretation. Although our findings are based on specific participant accounts and chatlogs, it is important to view our framework as just one interpretive lens among many. Moreover, as relational dissonance partially operates outside conscious awareness, our methodology captures its observable manifestations, rather than the complete psychological process of relational dissonance. The value of this construct lies in illuminating the continuous micro-negotiations between explicit framings and implicit dynamics, which cannot be fully captured by HCI taxonomies that assume stable roles.
@@
Empirical Limitations. Our findings emerge from a specific population of qualitative researchers operating within the professional norms of knowledge production, agency, and methodological reflexivity. We cannot claim that the relational dynamics we document are universal or representative of the general population. The tensions between tool-use framing and relational engagement might be especially pronounced in our study population, which has professional stakes in interpreting qualitative data. As entirely different modes of relational engagement may arise during everyday interactions with AI systems, generalizing our findings to a general audience is not straightforward. Therefore, our empirical contribution should be understood as documenting how relational dissonance manifests in knowledge-work contexts where professional integrity and epistemic authority are at stake, rather than identifying universal dynamics of human-AI interaction. However, our proposal that relational dissonance may be useful for understanding human-AI interaction more broadly extends beyond what the empirical evidence in this paper demonstrates. Whether our construct applies similarly across different populations, settings, and tasks remains a question for future research.
@@ 7. Conclusion
Acknowledgements.
This work was supported by the Jenny and Antti Wihuri Foundation. We gratefully acknowledge the Aalto Center for Qualitative Research (Qual+) and the House of AI at Aalto University for providing the platform that enabled this research. CC further acknowledges the research environment provided by ELLIS Institute Finland.
@@ Appendix C Study Details
In this appendix, we provide key details related to the setup of our study, sharing
(1) the mapping between the human-readable participant identifiers used in the paper and the random identifiers used in the study itself (Section C.1), (2) the survey questions and item scales used in the pre-session survey (Section C.2), (3) the precise wording of the fictional qualitative-research task to be addressed by participants through human-AI interactions (Section C.3), (4) the precise versions of the AI systems used by our participants (Section C.4), (5) the images presented during the metaphor-elicitation exercise (Section C.5), and (6) the structure underlying the group discussions (Section C.6).
@@ C.1. Participant Identifiers
The mapping of participant numbers to the identifiers used in the source data is shown in Table 1.
@@ C.2. Survey Questions and Item Scales
In the following, we reproduce the questions participants answered in our pre-session survey, organized by their variable group in Figure 3, along with the scales on which they could provide their answer (see Section 3.1). In addition to these questions, we also asked participants to self-report their age, nationality, and gender as basic demographic information. All questions were optional. An anonymized CSV file of all answers (with nationalities censored to protect privacy) is provided in the supplementary material.
@@ C.5. Images Provided for Selection as Visual Metaphors
Figure 5. Visual metaphors used in the image-selection exercise.
Grid of 34 diverse images used as visual metaphors during the image-selection exercise. Images include: abstract patterns, animals, human representations, technology, various objects, nature scenes, other symbolic imagery, such as a road sign ”Please Break for Snakes”, a hand reflecting from a mirror, and a throne. Images are arranged in no particular order.